Given f(x) = –2x – 1, find f(a – 3).

Answer by rapaljer(4551)

f(x) = –2x – 1
f(a–3)=-2(a-3) - 1
f(a-3)= -2a+6-1
f(a-3) = -2a+5

R^2 at SCC